julio has 5 more dollars than dan. altogether julio and dan have 19 dollars. how much mney does each young man have?
Answer by Jolliano(16)

Let J be the amount of money that Julio has.
Let D be the amount of money that Dan has.

J + D = 19 ....(1)
J - D = 5 ....(2)

Adding both equations we get,

2J = 24

J = 24/2

J = "12 Dollars"

Substitute the value of J in any equation above (I'll use Equation 1)

J + D = 19
12 + D = 19
D = 19 - 12
D = "7 Dollars".

Therefore, Julio has "12 Dollars" while Dan has "7 Dollars".